The following table is from the publication of Vital and Health Statistics from the Centers for Disease Control and Prevention/National Center for Health Statistics. The individuals in the following table have an eye irritation, a nose irritation, or a throat irritation. They have only one of the three. Is there sufficient evidence to reject the hypothesis that the type of ear, nose, or throat irritation is independent of the age group at a level of significance equal to .05?Age (years)

Type of Irritation

18-29 30-44 45-64 65 and Older

Eye 460 567 358 60

Nose 904 1311 793 109

Throat 236 322 159 17

Find Test Statistic

Find P-value 2
Answer #1

Here chi square test of independence will be used

Hypotheses are:

H0: The type of ear, nose, or throat irritation is independent of the age group.

Ha: The type of ear, nose, or throat irritation is not independent of the age group.

Following table shows the row total and column total:

18-29 30-44 45-64 65 and older Total
Eye 460 567 358 60 1445
Nose 904 1311 793 109 3117
Throat 236 322 159 17 734
Total 1600 2200 1310 186 5296

Expected frequencies will be calculated as follows:

E_{ij}=\frac{ith\,row\,total\cdot jth\,row\,total}{grand\,total}

Following table shows the expected frequency:

18-29 30-44 45-64 65 and older Total
Eye 436.556 600.264 357.43 50.75 1445
Nose 941.692 1294.826 771.01 109.472 3117
Throat 221.752 304.909 181.56 25.779 734
Total 1600 2200 1310 186 5296

Following table shows the calculations for test statistics:

O E (O-E)^2/E
460 436.556 1.25899343
904 941.692 1.508653428
236 221.752 0.915461885
567 600.264 1.843345088
1311 1294.826 0.202033537
322 304.909 0.957998226
358 357.43 0.000908989
793 771.01 0.627177469
159 181.56 2.80322538
60 50.75 1.685960591
109 109.472 0.002035077
17 25.779 2.989675356
Total 14.79546846

The test statistics

\chi^{2}=\sum \left ( \frac{(O-E)^{2}}{E} \right )=14.795

Degree of freedom: df =( number of rows -1)*(number of columns-1) = (3-1)*(4-1)=6

The p-value is:

p-value = 0.0219

Since p-value is less than 0.05 so we reject the null hypothesis.

There is sufficient evidence to reject the hypothesis that the type of ear, nose, or throat irritation is independent of the age group at a level of significance equal to 0.05.